MOA-II Obsevation Log Monthly Summary 2007/6
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 7083-8059 | 977 | 5301 | 300960 |
LMC | 1556-1580 | 25 | 156 | 33500 |
SMC | 385-392 | 8 | 16 | 3800 |
ALERT | 346-357 | 12 | 12 | 720 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 1022 | 5485 | 338980 (94h09m40s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
183 | 207 | 179 | 177 | 841 | 167 | 200 | 171 | 831 | 232 | 176 | 170 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
165 | 210 | 158 | 164 | 259 | 215 | 156 | 151 | 150 | 139 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
9 | 9 | 9 | 9 | 9 | 11 | 15 | 15 | 15 | 12 | 10 | 10 |
Field13 | Field14 | Field15 | Field16 |
11 | 12 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
8 | 8 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
Category | Target | Nrun |
ALERT | MOA2007BLG268 | 12 |
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/6/1: D2324(200sec) , D2325(20sec), D2330(30sec), D2331(40sec) and D2332(60sec) are OK. I changed exposure time frequently for GB observations because the moon is full. / night time = 46020(sec)
2007/6/2: GB is coverd by the moon. D2333(20sec) and D2334(100sec) are OK. / night time = 46080(sec)
2007/6/3: D2335(30sec), D2336(40sec), D2337(60sec) and D2338(200sec) are OK. / night time = 46080(sec)
2007/6/4: no observation / night time = 46200(sec)
2007/6/5: No obserbation / night time = 46260(sec)
2007/6/6: No observation / night time = 46260(sec)
2007/6/7: No observation / night time = 46320(sec)
2007/6/8: D2339-60-a,b,c is ok D2340-300-a,b,c is ok Observed gb17 every twe times for MOA189 / night time = 46380(sec)
2007/6/9: D2341-60-a,b,c is ok D2342-300-a,b,c is ok / night time = 46440(sec)
2007/6/10: D2343-60-a,b,c is ok D2344-300-a,b,c is ok / night time = 46440(sec)
2007/6/11: D2345 (60sec) , D2346 (300sec) is ok Pumping / night time = 46500(sec)
2007/6/12: D2347(300sec) ,D2348 (60sec) is ok / night time = 46500(sec)
2007/6/13: Do not use F1752-F1757 because of lens condencation , No observation / night time = 46560(sec)
2007/6/14: D2349 (60sec) is ok / night time = 46560(sec)
2007/6/15: No observation. / night time = 46560(sec)
2007/6/16: observed gb7 as often as gb5,9 for MOA212 alerted by PLANET,and gb10 D2350(60sec) is ok / night time = 46620(sec)
2007/6/17: D2351(60sec),D2352(300sec) is ok / night time = 46620(sec)
2007/6/18: condensed on the lens2 shutter broken !! / night time = 46620(sec)
2007/6/19: D2359(60sec) is OK. Shutter fixed. camera mounted. We executed telescope analysis. chip2 bias value suddenly high ~2000ADU. Resistance between the camera and the telescope is about 9kohm, but CCD noise is not increased. / night time = 46680(sec)
2007/6/20: No observation,Took flat with many rotator angle ( F1800-1813 ) and for gain ( F1800-1805,1808-1809,1811-1818) ,Took dark for flat (D2365-2381) / night time = 46680(sec)
2007/6/21: No observation / night time = 46620(sec)
2007/6/22: Noi observation / night time = 46620(sec)
2007/6/23: dark D2386 (60sec) was OK. / night time = 46620(sec)
2007/6/24: sky is high ADU due to the Moon and cloud (~2:00),observed gb2 (for MOA211),10 (for MOA209,246),14 (for MOA217),18 (for MOA249) / night time = 46680(sec)
2007/6/25: D2391(60sec) ,D2392(200sec) are OK ,observed gb10 for MOA109 and MOA146 and gb18 for MOA249 as often as gb5,9 / night time = 46620(sec)
2007/6/26: We observed intensively gb18 for MOA249, gb10 for MOA209 and gb14 for MOA 217. Good seeing !! Dark D2399(200sec), D2400(60sec) are OK. / night time = 46620(sec)
2007/6/27: We took a special field for MOA268 as often as gb5,9. DARK D2405(60sec), D2406(200sec) are OK. We exchanged the broken vacuum-sensor (cold-cathode type) for the new one (hot-cathode type). / night time = 46620(sec)
2007/6/28: No observation. / night time = 46560(sec)
2007/6/29: No observation. / night time = 46560(sec)
2007/6/30: DARK D2413(150sec) is OK, We couldn't observe the GB because it was very close to the full moon. We've done pumping for a short time. / night time = 46560(sec)
TOTAL NIGHT TIME: 1394460(sec)
OPERATING RATE: 40.0 %
GB: 36.8 %
LMC: 2.8 %
SMC: 0.3 %
ALERT: 0.1 %
GRB: 0.0 %
Return to index
Previous month
Next month